OpenStudy (anonymous):

perform the division 19x^7+6x^3/x

OpenStudy (anonymous):

@campbell_st u take over

OpenStudy (campbell_st):

write it as the sum of 2 fractions \[\frac{19x^7}{x} + \frac{6x^3}{x}\] index law for division is subtract the powers \[19x^{7-1} + 6x^{3-1} = 19x^6 + 6x^2\]

OpenStudy (anonymous):

19x^6+6x^2

OpenStudy (anonymous):

is that how

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

@bh104 camp is correct